# Action-OOD: An End-to-End Skeleton-Based Model for Robust Out-of-Distribution Human Action Detection

### v1

@@ Abstract
Human action recognition is a crucial task in computer vision systems. However, in real-world scenarios, human actions often fall outside the distribution of training data, requiring a model to both recognize in-distribution (ID) actions and reject out-of-distribution (OOD) ones. Despite its importance, there has been limited research on OOD detection in human actions. Existing works on OOD detection mainly focus on image data with RGB structure, and many methods are post-hoc in nature. While these methods are convenient and computationally efficient, they often lack sufficient accuracy and fail to consider the presence of OOD samples. To address these challenges, we propose a novel end-to-end skeleton-based model called Action-OOD, specifically designed for OOD human action detection. Unlike some existing approaches that may require prior knowledge of existing OOD data distribution, our model solely utilizes in-distribution (ID) data during the training stage, effectively mitigating the overconfidence issue prevalent in OOD detection. We introduce an attention-based feature fusion block, which enhances the model’s capability to recognize unknown classes while preserving classification accuracy for known classes. Further, we present a novel energy-based loss function and successfully integrate it with the traditional cross-entropy loss to maximize the separation of data distributions between ID and OOD. Through extensive experiments conducted on NTU-RGB+D 60, NTU-RGB+D 120, and Kinetics-400 datasets, we demonstrate the superior performance of our proposed approach compared to state-of-the-art methods. Our findings underscore the effectiveness of classic OOD detection techniques in the context of skeleton-based action recognition tasks, offering promising avenues for future research in this field. Code will be available at: https://github.com/YilliaJing/Action-OOD.git.
@@ 1 Introduction
Human-centric visual understanding enhances the development of intelligent machines, enabling humans to accomplish greater tasks and improve their quality of life. Deep learning-based models have shown remarkable performance in supervised human action classification tasks. In skeleton-based human action datasets, this task can be solved as a typical spatio-temporal modeling problem (Yan et al., 2018; Chen et al., 2021; Cheng et al., 2020). Although these methods have achieved high accuracy in supervised classification problems, they still lack consideration for out-of-distribution (OOD) issues. An important assumption in these approaches is that the distribution of the training set is similar to that of the test set. However, this can not always be achieved in practice. When these models are applied to recognize samples outside the distribution of the training set, they may exhibit overconfidence by selecting a category from the known distribution (Yu et al., 2024). Hence, it is essential for such models to be aware of the boundaries of learning for safety concerns (Nalisnick et al., 2019), ensuring the robustness of computer vision systems. How to deal with this OOD detection problem has attracted widespread attention in fields such as autonomous driving (Geiger et al., 2012) and medical image analysis (Schlegl et al., 2017).
@@
Works in OOD detection focus on OOD samples caused by semantic shift, aiming to determine whether a sample belongs to a category included in the training dataset. This task presents two challenges: quantifying the differences between in-distribution (ID) and OOD samples, and distinguishing them based on their expressions. For the first challenge, one typical approach is score measuring methods (Hendrycks and Gimpel, 2017), such as the softmax confidence score (Hendrycks and Gimpel, 2017). Inputs with low softmax confidence scores are classified as OOD. However, neural networks can produce arbitrarily high softmax confidence scores for inputs far from the training data (Nguyen et al., 2015). To address this, the energy score is proposed for OOD detection. It can be derived from a purely discriminative classification model without explicitly relying on a density estimator. Experiments show that it outperforms softmax-based score measuring methods.
@@
Based on such score measuring methods, a crucial issue is how to maximize the difference in score distributions between OOD and ID samples. To achieve this, difference amplification methods (Sun et al., 2021; Zhu et al., 2022) have been designed. However, these methods are mostly implemented in a post-hoc manner, meaning the model itself cannot inherently distinguish OOD samples. This does not fundamentally solve the problem of ’letting the model know what it knows.’ Furthermore, these methods were originally designed for OOD detection in RGB-structured datasets. Their effectiveness in handling action recognition problems with skeleton structure data has not yet been investigated.
@@
Furthermore, methods by Wu and et al. (2023) have attempted to achieve this goal by introducing OOD distribution information into the training stage. However, in practice, we cannot measure the specific categories of OOD. Similarly, when the OOD category itself expands, the overconfidence problem still remains. In summary, while these methods significantly improve the accuracy of OOD recognition, achieving this goal using an end-to-end framework without prior OOD information remains challenging.
@@
To address the limits and challenges mentioned above, we propose an end-to-end model called Action-OOD to deal with the out-of-distribution detection problem for skeleton-based action recognition. Considering that the state-of-the-art feature extractors for supervised skeleton-based recognition are well-developed, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for the supervised classes. We use HD-GCN (Lee et al., 2023), a prevailing backbone at present, as our main feature extractor. Based on its basic output, we first creatively applied a feature activation module that utilizes ASH (Djurisic and et al., 2023) to project the original features for fitting OOD detection in the training stage. Specifically, it filters useless information and amplifies the key feature dimensions to generate a more sensitive and accurate representation of abnormal detections. Then, regarding this as an orthogonal aggregation to the original features, we innovatively designed a feature fusion module, which integrates the original backbone outputs and the OOD-targeted activated features to enable the model to preserve a strong recognition ability for ID data while also nicely discriminating the potential OOD existences. To support the framework with end-to-end training, we suggested a meticulously designed learning loss that combines the energy function and cross-entropy information. This approach aims to improve the model’s understanding of the energy score distribution in ID data, thereby further benefiting the accurate detection of OOD samples.
@@
    We propose Action-OOD, an end-to-end skeleton-based OOD detection model that addresses overconfidence by solely training on ID data.
@@
    The proposed model Action-OOD outperforms baselines on NTU-RGB+D 60, NTU-RGB+D 120, and Kinetics-400 datasets in the human action OOD detection task.
The rest of the paper is organized as follows: Section 2 explains basic knowledge of skeleton-based action recognition tasks with GCNs and the theory of why energy score is most widely used in OOD detection. Section 3 introduces our proposed method in detail, including proposed framework, feature fusion block and loss function design. Section 4 mainly discusses the result of experiments. We offer conclusion and discuss future works in Section 5.
@@ 2.1 GCNs for Skeleton based Action Recognition
Human action recognition task is a typical problem in computer vision, which has wide applications in the real world including human-computer interaction (Nikam and Ambekar, 2016) and video surveillance (Jiang et al., 2015). Skeleton-based method has its unique advantages. The skeleton structure can provide more accurate node-level information and ignore the influence of background noise information, thus improving the prediction accuracy of recognition.
In the early phases of skeleton-based action recognition using deep learning methods, convolution neural networks (CNNs) (Chéron et al., 2015; Liu et al., 2017b; Simonyan and Zisserman, 2014), and recurrent neural networks (RNNs) (Wang and Wang, 2017; Liu et al., 2017a) were commonly utilized. However, these approaches had limitations as they didn’t fully leverage the structural arrangement of the joints.
After that, researchers build spatio-temporal architechers to handle this problem. Many works use handcrafted physically connected (PC) edges among human skeletons to extract spatial features (Chen et al., 2021; Cheng et al., 2020) and GCNs can handle this type of graph-structured data very well. The graph defined on the human skeleton is denoted as $G(V,E)$, with $V$ representing the joint group and $E$ representing the edge group. The representation of 3D time-series skeletal data is denoted as $X\in\mathbb{R}^{C\times T\times|V|}$, where $|V|$ represents the number of joint nodes, $C$ is the number of channel, and $T$ denotes the temporal window size. The operation of GCN with an input feature map $X$ can be described as follows:
@@
Here, $\boldsymbol{\Theta}$ denotes the pointwise convolution operation. $A$ represents adjacency matrix that is initialized as $\mathbf{\Lambda}^{-\frac{1}{2}}\mathbf{A}\boldsymbol{\Lambda}^{-\frac{1}{2}}\in\mathbb{R}^{N_{g}\times|V|\times|V|}$ , where $\Lambda$ is a diagonal matrix for normalization and $N_{g}=3$ in experiments.
However, relying solely on PC edges for spatial relationships is insufficient, as it limits the receptive fields due to their heuristic and fixed nature. Additionally, the varying contributions of each edge must be considered, as the importance of joints in different body parts varies for specific actions. To address these issues, Hierarchically Decomposed Graph Convolutional Networks (HD-GCN) (Lee et al., 2023) were proposed. The graph convolution operation of HD-GCN is defined as:
@@
where $G=\left\{{g_{pc},g_{sl},g_{fc}}\right\}$ denotes three graph subsets, and $g_{pc}$, $g_{sl}$, and $g_{fc}$ indicate physical connections, self-loops, and fully connected joint subsets, respectively. It represents three typical useful graph relationships to help explore spatial relationships between joints. The construction of these three kinds will be introduced in detail in section 3.1. With this operation, HD-GCN achieved an average Top-1 accuracy of 93.9% across three experimental datasets.
@@ 2.2 Methods for out-of-distribution detection
Research on out-of-distribution detection can be mainly categorized into two ways. First, we can define a scoring function that maps each input point to a single scalar, such that in-distribution and out-of-distribution data will have different distributions. Thus, we can identify the out-of-distribution samples by calculating this score. There are many works about how to define this kind of score for pre-trained neural networks. Hendrycks and Gimpel (2017) proposed the maximum predicted softmax probability (MSP). After that, Wu and et al. (2023) proved that the log of the MSP was equivalent to a special case of the free energy score. This indicates that there are some cases in ID and OOD data with similar MSP values but different energy scores. As a result, energy score has become one of the most widely used score measuring method.
@@
Secondly, we can distinguish ID and OOD by trying to modify the feature captured by a pre-trained neural networks in a post-hoc way, which is usually applied to feature activations. Numerous experiments have shown that different feature activation methods can help differentiate ID and OOD. Sun et al. (2021) proposed rectified activation (ReAct) after observing that OOD data can trigger unit activation patterns that were significantly different from ID data in the penultimate layer of the model. Performing truncation can help drastically improve the separation of ID and OOD data. Zhu et al. (2022) found that some operations similar to Batch Normalization might increase the difference in scores (e.g., MSP and energy score) between ID and OOD data. DICE (Sun and Li, 2022) was proposed to address the shortcomings of redundant information expressed in the high space of neural networks. It attempted to define weights (i.e., weight × activation) to sort and filter the nodes in the penultimate layer of the neural network for denoising. This further reduces the variance value of the ID and OOD data score distributions, making it easier to separate the peaks of the two distributions. Similar to the previous works, Djurisic and et al. (2023) removed the top-K elements (usually a large portion over 50%) and adjust the remaining (e.g. 10%) activation values by scaling them up or straightly assign them a constant number. It achieves state-of-art performance on several benchmarks of image classification.
@@
However, most of the first two kinds of methods are considered post-hoc. While they can reduce calculation costs than training another new specific OOD detection model, research by Liu et al. (2023) showed that they were still less competitive than some works that utilized OOD information in the training stage, such as works by Hendrycks et al. (2022) and Song et al. (2022). Additionally, these methods were initially designed for OOD detection in RGB-structured datasets. Their effectiveness in addressing action recognition problems with skeleton structure data has not yet been explored. In this work, we prove the validity of these methods on skeleton-based datasets and then take steps to solve these problems.
Figure 1: Framework of the proposed Action-OOD model. The symbol $\oplus$ represents the vector concatenate operation.
@@ 2.3 Energy-based out-of-distribution detection
In deep neural networks, out-of-distribution detection distinguishes samples that deviate from the training distribution. In this work, we only focus on the standard OOD detection that concerns semantic shifts, in which OOD data is defined as test samples from semantic categories that were not seen during training. Ideally, the neural network should be able to reject such samples as OOD while maintaining strong classification performance on ID test samples belonging to seen training categories (Zhang and et al., 2023). To achieve this goal, a common approach is to utilize the data density function, denoted as $p^{in}(x)$, and identify instances with low likelihood as OOD. Nevertheless, prior studies have demonstrated that density functions estimated by deep generative models are not consistently reliable for OOD detection (Nalisnick et al., 2019).
Energy-based out-of-distribution detection employs an energy score for detection, where the difference of energies between ID and OOD facilitates separation. The energy score addresses a crucial issue associated with softmax confidence, which can lead to excessively high values for OOD examples (Hein et al., 2019).
@@
$E(\mathbf{x};f)=-T\cdot\log\sum_{i}^{K}e^{f_{i}(\mathbf{x})/T},$ (3)
where $T$ is the temperature parameter. It can also be proved that a model trained with negative log-likelihood (NLL) loss will push down energy for in-distribution data points and pull up for other labels (LeCun et al., 2006). Since cross-entropy loss can be implemented as NLL loss after the logSoftmax function, it will also push down the energy score of ID data. In other words, the ID sample will have a larger negative energy score $-E(x;f)$ than the OOD sample.
@@ 3 Method
Our main objective is to achieve accurate human action OOD detection through an end-to-end framework without any other subsequent fine-tuning or post-hoc operations. The whole framework of the proposed skeleton-based OOD human action detection model is shown in Fig. 1.
It can be divided into three main steps of the end-to-end Action-OOD framework: Firstly, we use HD-GCN (Lee et al., 2023), a GCN framework that achieves state-of-the-art accuracy in ID classification, to perform feature extraction on skeleton data. Secondly, The extracted features are then activated and fused with the original ones. The function of the feature fusion block is to improve the recognition ability of OOD samples while preserving the original classification performance of ID data. Finally, it is fed into the classifier, and OOD detection is performed based on the classifier output logits as shown in Fig. 4. Besides, the whole model is trained by an energy-based loss function. Our work focuses on feature processing and OOD detection. Below, we will delve extensively into each of these components.

### v2

# Skeleton-OOD: An End-to-End Skeleton-Based Model for Robust Out-of-Distribution Human Action Detection

@@ Abstract
Human action recognition is crucial in computer vision systems. However, in real-world scenarios, human actions often fall outside the distribution of training data, requiring a model to both recognize in-distribution (ID) actions and reject out-of-distribution (OOD) ones. Despite its importance, there has been limited research on OOD detection in human actions. Existing works on OOD detection mainly focus on image data with RGB structure, and many methods are post-hoc in nature. While these methods are convenient and computationally efficient, they often lack sufficient accuracy, fail to consider the exposure of OOD samples, and ignore the application in skeleton structure data. To address these challenges, we propose a novel end-to-end skeleton-based model called Skeleton-OOD, which is committed to improving the effectiveness of OOD tasks while ensuring the accuracy of ID recognition. Through extensive experiments conducted on NTU-RGB+D 60, NTU-RGB+D 120, and Kinetics-400 datasets, Skeleton-OOD demonstrates the superior performance of our proposed approach compared to state-of-the-art methods. Our findings underscore the effectiveness of classic OOD detection techniques in the context of skeleton-based action recognition tasks, offering promising avenues for future research in this field. Code is available at https://github.com/YilliaJing/Skeleton-OOD.git.
@@ 1 Introduction
Human-centric visual understanding enhances the development of intelligent machines, enabling humans to accomplish greater tasks and improve their quality of life. Deep learning-based models have shown remarkable performance in supervised human action classification tasks. In skeleton-based human action datasets, this task can be solved as a typical spatio-temporal modeling problem [1, 2, 3]. Although these methods have achieved high accuracy in supervised classification problems, they still lack consideration for out-of-distribution (OOD) issues. An important assumption in these approaches is that the distribution of the training set is similar to that of the test set. However, this can not always be achieved in practice. When straightly applying these action recognition models to recognize samples outside the distribution of the train dataset, they may exhibit overconfidence by selecting a category from the known distribution [4]. Hence, it is essential for such models to be aware of the boundaries of learning for safety concerns [5], ensuring the robustness of computer vision systems. How to deal with this OOD detection problem has attracted widespread attention in fields such as autonomous driving [6] and medical image analysis [7].
@@
Works in OOD detection focus on OOD samples caused by semantic shift, aiming to determine whether a sample belongs to a category included in the training dataset. This task presents two challenges: quantifying the differences between in-distribution (ID) and OOD samples and distinguishing them based on their expressions. For the first challenge, one typical approach is score measuring methods [8], such as the softmax confidence score [8]. Inputs with low softmax confidence scores are classified as OOD. However, neural networks can produce arbitrarily high softmax confidence scores for inputs far from the training data [9]. To address this, the energy score is proposed for OOD detection. It can be derived from a purely discriminative classification model without explicitly relying on a density estimator. Experiments have shown that it outperforms softmax-based score measuring methods.
@@
Based on such score measuring methods, a crucial issue is how to maximize the difference in score distributions between OOD and ID samples. To achieve this, difference amplification methods [10, 11] have been designed. However, these methods are mostly implemented in a post-hoc manner, meaning the model itself cannot inherently distinguish OOD samples. This does not fundamentally solve the problem of ‘letting the model know what it knows’. Furthermore, these methods were originally designed for OOD detection in RGB-structured datasets. Their effectiveness in handling action recognition problems with skeleton structure data has not yet been investigated.
@@
Furthermore, methods like [12, 13] have attempted to achieve this goal by introducing OOD related information into the training stage. However, in practice, we cannot measure the specific categories of OOD. Similarly, when the OOD category itself expands, the overconfidence problem still remains. In summary, while these methods significantly improve the accuracy of OOD recognition, achieving this goal using an end-to-end framework without prior OOD information remains challenging.
@@
To address the limits and challenges mentioned above, we propose an end-to-end model called Skeleton-OOD to deal with the out-of-distribution detection problem for skeleton-based action recognition. Considering that the state-of-the-art feature extractors for supervised ID classification are well-developed, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for the ID classification. We input graphs, joints, and temporal sequence into the spatial-temporal feature extractor to obtain feature embeddings as output. Based on this, we first creatively applied a feature activation module that utilizes ASH [14] to project the original features for fitting OOD detection in the training stage. Specifically, it filters useless information and amplifies the key feature dimensions to generate a more sensitive and accurate representation of abnormal detections. However, in its original work, the activation method was only used in a post hoc manner after the trained model, and the energy score was only used for judgment. Then, regarding this as an orthogonal aggregation to the original features, we innovatively designed a feature fusion module, which integrates the original backbone outputs and the OOD-targeted activated features to enable the model to preserve a strong recognition ability for ID data while also nicely discriminating the potential OOD existences. To support the framework with end-to-end training, we suggested a meticulously designed learning loss that combines the energy function and cross-entropy information. This approach aims to improve the model’s understanding of the energy score distribution in ID data, thereby further benefiting the accurate detection of OOD samples.
@@
    We propose Skeleton-OOD, an end-to-end skeleton-based OOD detection model that addresses overconfidence by solely training on ID data.
@@
    The proposed model Skeleton-OOD outperforms baselines on NTU-RGB+D 60, NTU-RGB+D 120, and Kinetics-400 datasets in the human action OOD detection task.
The rest of the paper is organized as follows: Section 2 explains basic knowledge of skeleton-based action recognition tasks with GCNs and the theory of why energy score is most widely used in OOD detection. Section 3 introduces our proposed method in detail, including the proposed framework, feature fusion block, and loss function design. Section 4 mainly discusses the results of experiments and thus shows the effectiveness and usability of our proposed model. We offer conclusions, limitations, and future works in Section 5.
@@ 2.1 GCNs for Skeleton based Action Recognition
Human action recognition tasks are a typical problem in computer vision, which have wide applications in the real world, including human-computer interaction [15] and video surveillance [16]. Skeleton-based methods have their unique advantages. The skeleton structure can provide more accurate node-level information and ignore the influence of background noise, thus improving the prediction accuracy of recognition.
In the early phases of skeleton-based action recognition using deep learning methods, convolutional neural networks (CNNs) [17, 18, 19], and recurrent neural networks (RNNs) [20, 21] were commonly utilized. However, these approaches had limitations as they didn’t fully leverage the structural arrangement of the joints.
Subsequently, researchers built spatio-temporal architectures to handle this problem. Many works use handcrafted physically connected (PC) edges among human skeletons to extract spatial features [2, 3], and GCNs can handle this type of graph-structured data very well. The graph defined on the human skeleton is denoted as $G(V,E)$, with $V$ representing the joint group and $E$ representing the edge group. The representation of 3D time-series skeletal data is denoted as $X\in\mathbb{R}^{c\times t\times|V|}$, where $|V|$ represents the number of joint nodes, $c$ is the number of channels, and $t$ denotes the temporal window size. The operation of GCN with an input feature map $X$ can be described as follows:
@@
Here, $\boldsymbol{\Theta}$ denotes the pointwise convolution operation. The adjacency matrix $A$, which is initialized as $\mathbf{\Lambda}^{-\frac{1}{2}}\mathbf{A}\boldsymbol{\Lambda}^{-\frac{1}{2}}\in\mathbb{R}^{N_{g}\times|V|\times|V|}$, where $\Lambda$ is a diagonal matrix for normalization, and $N_{g}=3$ in experiments.
However, relying solely on PC edges for spatial relationships is insufficient, as it limits the receptive fields due to their heuristic and fixed nature [22]. Additionally, the varying contributions of each edge must be considered, as the importance of joints in different body parts varies for specific actions. To address these issues, Hierarchically Decomposed Graph Convolutional Networks (HD-GCN) [23] were proposed. The graph convolution operation of HD-GCN is defined as:
@@
where $G=\left\{{g_{pc},g_{sl},g_{fc}}\right\}$ denotes three graph subsets, and $g_{pc}$, $g_{sl}$, and $g_{fc}$ indicate physical connections, self-loops, and fully connected joint subsets, respectively. It represents three typical useful graph relationships to help explore spatial relationships between joints. The construction of these three types of subsets will be introduced in detail in section 3.1. With this operation, HD-GCN achieved an average Top-1 accuracy of 93.9% across three experimental datasets.
@@ 2.2 Methods for out-of-distribution detection
Research on out-of-distribution detection can be mainly categorized into two ways. First, we can define a scoring function that maps each input point to a single scalar, such that in-distribution and out-of-distribution data will have different distributions. Thus, we can identify the out-of-distribution samples by calculating this score. There are many works about how to define this kind of score for pre-trained neural networks. [8] proposed the maximum predicted softmax probability (MSP). After that, [12] proved that the log of the MSP was equivalent to a special case of the free energy score. This indicates that there are some cases in ID and OOD data with similar MSP values but different energy scores. As a result, energy score has become one of the most widely used score measuring method.
@@
Secondly, we can distinguish ID and OOD by trying to modify the feature captured by a pre-trained neural network in a post-hoc way, which is usually applied to feature activations. Numerous experiments have shown that different feature activation methods can help differentiate ID and OOD. [10] proposed rectified activation (ReAct) after observing that OOD data can trigger unit activation patterns that were significantly different from ID data in the penultimate layer of the model. Performing truncation can help drastically improve the separation of ID and OOD data. [11] found that some operations similar to Batch Normalization might increase the difference in scores (e.g., MSP and energy score) between ID and OOD data. DICE [24] was proposed to address the shortcomings of redundant information expressed in the high space of neural networks. It attempted to define weights (i.e., weight × activation) to sort and filter the nodes in the penultimate layer of the neural network for denoising. This further reduces the variance value of the ID and OOD data score distributions, making it easier to separate the peaks of the two distributions. Similar to the previous works, [14] removed the top-K elements (usually a large portion over 50%) and adjusted the remaining (e.g. 10%) activation values by scaling them up or straightly assigned them a constant number. It achieves state-of-art performance on several benchmarks of image classification.
@@
However, most of the first two kinds of methods are considered post-hoc. While they can reduce calculation costs than training another new specific OOD detection model, research by [25] showed that they were still less competitive than some works that utilized OOD information in the training stage, such as works by [26] and [27]. Additionally, these methods were initially designed for OOD detection in RGB-structured datasets. Their effectiveness in addressing action recognition problems with skeleton structure data has not yet been explored. In this work, we prove the validity of these methods on skeleton-based datasets and then take steps to solve these problems.
@@ 2.3 Energy-based out-of-distribution detection
In deep neural networks, out-of-distribution detection distinguishes samples that deviate from the training distribution. In this work, we only focus on the standard OOD detection that concerns semantic shifts, in which OOD data is defined as test samples from semantic categories that were not seen during training. Ideally, the neural network should be able to reject such samples as OOD while maintaining strong classification performance on ID test samples belonging to seen training categories [28]. To achieve this goal, a common approach is to utilize the data density function, denoted as $p^{in}(x)$, and identify instances with low likelihood as OOD. Nevertheless, prior studies have demonstrated that density functions estimated by deep generative models are not consistently reliable for OOD detection [5].
Energy-based out-of-distribution detection employs the energy score for detection, where the difference of energies between ID and OOD facilitates separation. The energy score addresses a crucial issue associated with softmax confidence, which can lead to excessively high values for OOD examples [29].
@@
$E(\mathbf{x};f)=-\varepsilon\cdot\log\sum_{i}^{K}e^{f_{i}(\mathbf{x})/\varepsilon},$ (3)
where $\varepsilon$ is the temperature parameter. It can also be proved that a model trained with negative log-likelihood (NLL) loss will push down energy for in-distribution data points and pull up for other labels [30]. Since cross-entropy loss can be implemented as NLL loss after the logSoftmax function, it will also push down the energy score of ID data. In other words, the ID sample will have a larger negative energy score $-E(x;f)$ than the OOD sample.
[…]
@@ 3 Method
Our goal is to achieve accurate human action OOD detection through an end-to-end framework without any other subsequent fine-tuning or post-hoc operations. The whole framework of the proposed skeleton-based OOD human action detection model is shown in Figure 1.
It can be divided into three main steps of the end-to-end Skeleton-OOD framework: Firstly, a GCN-based backbone is employed to extract features from skeleton data. Secondly, The extracted features are then activated and fused with the original ones. The function of the feature fusion block is to improve the recognition ability of OOD samples while preserving the original classification performance of ID data. Finally, it is fed into the classifier, and OOD detection is performed based on the classifier output logits as shown in Figure 4. Besides, the whole model is trained by our newly designed energy-based loss function. In the following sections, we will provide a detailed discussion of each component.